# Deanonymizing Ethereum Validators: The P2P Network Has a Privacy Issue

### v1

@@ 7 Insights
FR SO VA ZH
FR 14,199 2,840 5,284 10,580
SO 2,840 11,155 5,227 9,215
VA 5,284 5,227 47,269 25,898
ZH 10,580 9,215 25,898 129,521
Table 6: Overlap of deanonymized validators across machines. With the exception of the ZH node, half of the validators deanonymized on a Rainbow node have also been deanonymized by at least one other Rainbow node.
@@ 7.1 Geographical Distribution
On a continent basis, 46.67% of the peers we deanonymize are located in Europe, 38.14% in North America, 9.98% in Asia, 4.17% in Oceania, 0.77% in South America, and 0.27% in Africa. In Figure 6(a), we show the number of peers on which we deanonymized validators by Rainbow node as well as overall per country for the top ten countries. We deanonymized the biggest number of peers in the United States making up 33.11% of the peers. This is more than double the number of peers we deanonymize in the next two biggest countries. This is similar to the overall geographical distribution of peers in the network, where the United States and Germany make up the biggest proportions. Additionally, we also observe a geographical bias of our Rainbow nodes in terms of the peers on which they deanonymize. Most notably, the VA Rainbow node (located in the United States) deanonymizes more peers in the United States than the ZH Rainbow node (located in central Europe). Similarly, the ZH node deanonymizes more peers in most European countries in comparison to the VA node.
@@
In Figure 6(b) we further provide insights regarding the number of validators deanonymized by each Rainbow node per country. We start by highlighting the most striking difference to the previous figure: most validators we locate are hosted in the Netherlands (i.e., 14.08%), while the United States only comes in fourth place. Note that a difference in the figures is expected given a non-uniform distribution of validators across machines. In general, we observe that there is a larger proportion of validators in Europe as opposed to peers with validators. In particular, 71.53% of the validators we locate are in Europe, 11.95% are in North America, 11.52% are in Asia, 4.90% are in Oceania, 0.06% are in Africa and 0.03% are in South America. Additionally, we again notice the same geographical bias, i.e., the SO node’s high relative proportion of deanonymizations in Australia and South Korea.
@@
We further repeat the preceding analysis for organizations instead of countries in Figure 7. The largest organization in terms of the number of peers is Comcast a residential ISP with 5.99% and the following three (Hetzner, OVH, and Amazon) are cloud providers (see Figure 7(a)). Around half of the peers we deanonymize are hosted through cloud providers, whereas the other half run through residential ISPs, i.e., home stakers.
@@
There is a large shift in these figures when we turn from peers (see Figure 7(a)) to validators (see Figure 7(b)). For one, eight out of the ten largest organizations are cloud providers and all of the biggest seven are cloud providers. We locate the largest number of validators in Amazon data centers, i.e., 19.80% of the validators we locate. Surprisingly, the ZH node deanonymizes by far the largest proportion of validators in Amazon data centers even though it is the only Rainbow node not run from an Amazon data center. Finally, we highlight that around 90% of the validators are run through cloud providers, with the other 10% belonging to residential ISPs.
@@ 7.2 Staking Pools
In Figure 8, we plot a cdf of the number of validators hosted per peer for these five staking pools. Recall in our earlier analysis (see Figure 4) the observation that node operators of staking pools tend to run a “round” number of validators on the same machine. We observe these same increases in Figure 8. Here we observe an average of 678 validators on a given peer, with the larger validator set containing 19,263 validators. We highlight the security concern this raises: though there are a million validators, just hundreds of these peers going offline can stall Ethereum as more than two-thirds of the validators must be online to guarantee liveness.
@@
8 Implications
The feasibility of deanonymizing a large chunk of Ethereum’s validators in the P2P network we demonstrate stands in opposition to its design principles. In the following, we outline the implications and security concerns stemming from this lack of privacy in the Ethereum P2P network, as well as insights into the distributions of validators across peers. We commence with the former.
Taking-out Preceding Block Proposers. A further security concern that has been discussed online is that there are incentives for the proposer of slot $n+1$ to prevent the proposer of slot $n$ from making a block proposal [17]. In doing so the proposer of slot $n+1$ would be expected to receive higher execution layer rewards, i.e., fees paid by transactions for (preferential) block inclusion, as the proposer could place both the transactions that would have been included by the previous proposers as well as any that arrive in the meantime in their block. The attacker (i.e., the block proposer of slot $n+1$) knows the ID of their victim (the block proposer of slot $n$) in the consensus layer in advance, as the proposer assigned is known at least one epoch (approximately six minutes) in advance. Now if the attacker deanonymizes the victim in the P2P network, the attacker could try to temporarily break the connection between the victim and the P2P network through a (D)DoS attack or BGP hijacking. Note that the attack only needs to be sustained for four seconds, as a block proposer has a four-second window to make their proposal.
Breaking Liveness and Safety. An attacker could extend the previous attack to break the liveness or safety of Ethereum. To break liveness, an attacker could continuously perform a DoS attack on the upcoming block proposer. If the attacker is successful in preventing the propagation of more than one-third of blocks, liveness comes to a halt as the finality gadget will not be able to collect a large enough quorum to finalize blocks. Alarmingly, an attacker can perform an attack on the safety of the blockchain. Many light clients on Ethereum optimistically consider the head of the chain finalized. Thus, breaking the synchrony assumptions through targeted DoS attacks or network partitioning could lead to safety issues [19]. Such attacks are much more elaborate but benefit from our proposed deanonymization. Further, recent work demonstrated the general susceptibility of the liveness of consensus protocols to DoS attacks [35].
Breaking Core Assumptions of Danksharding. Importantly, the possibility of deanonymizing validators in the P2P network is in stark contrast with the assumptions required for danksharding – a central element of the Ethereum roadmap to scale the blockchain [14], targeting how much data can be stored by the network. Danksharding [55] plans to increase the block size by orders of magnitude (i.e., to 30 MB). As a result, validators will no longer be expected to store the entire chain but each validator will only store a small subset of a block’s data. In this setting, nodes still need to be able to reach agreement that all data is being stored and available, for which they use data-availability sampling (DAS). DAS ensures that the data for a block is available by having clients randomly sample batches [55, 53]. However, DAS requires source anonymity to prevent a data withholding attack in which a malicious validator convinces the network that the data is available while not serving it to clients that need it [1]. In the absence of source anonymity, which we show to not be fulfilled in the Ethereum P2P network, the adversary can serve requests selectively and break the probabilistic guarantees of random sampling that protect against such attacks [67].
@@ 8 Implications
Decentralization. Decentralization of the P2P network as well as Ethereum’s consensus layer is essential to prevent one party from maliciously overtaking the network. The decentralization of Ethereum’s PoS consensus layer is a topic of frequent discussion, given that the largest staking pool Lido controls nearly a third of the staking power – a critical safety threshold in Ethereum PoS. To combat the criticism, Lido proclaims that their validator set is operated by multiple independent node operators and thus they should all be seen as independent entities. While it is unclear whether this characterization is accurate [36], our insights into the P2P network reveal a further entanglement between different staking pools. As we outline in Section 7.2, not only do node operators run validators for different staking pools but we even find that they run them from the same nodes. This shows that staking pools are not always independent from each other, e.g., whenever such a node goes offline, validators from both protocols go offline. Similarly, we find that some peers host thousands of validators. Our insights gained from the deanonymization raise questions regarding the decentralization and resilience of the Ethereum consensus layer.
@@ 9.1 Providing Anonymity
We begin by discussing techniques that enhance anonymity within the P2P network. Before delving into the details, it is important to note that the current max effective balance for a validator is 32 ETH, which results in many redundant validators being run and controlled by the same entity. However, in the upcoming Pectra hardfork [56], the max effective balance is set to increase to 2048 ETH [51]. This change will allow entities to consolidate up to 64 validators into a single entity. Consequently, the next hard fork has the potential to reduce the number of attestations significantly propagated through the P2P network. This reduction allows for the following approaches that slightly increase the number of P2P messages to preserve anonymity. Importantly, with such an increase in P2P messages, the reductions in signature verification (i.e., one per validator per attestation) remain preserved.
@@
Solutions based on adding more nodes don’t tackle the root cause of the issue (likely just increasing the complexity of deanonymization) and almost universally lead to higher message complexity. Furthermore, they also result in larger operating costs, that are especially difficult to bear for solo stakers, thus going against the ethos of Ethereum.
Private Peering Agreements. Out of the box, the Lighthouse and Prysm Ethereum clients provide the capacity to perform private peering agreements – a set of trusted peers can be defined who function as additional relays for gossip messages. This functionality is meant for increased performance and reliability, however, it also provides k-anonymity***In this context, a node is said to be k-anonymous if it cannot be distinguished from k-1 other nodes.. The positive effect of such a mitigation can be seen in validators who make use of service providers at the gossip layer (see Section 6). Such practices make it harder to uniquely map validators to one IP, and one would need to target the $k$ peers in the peering agreement during an attack. Private peering agreements might be established off-band, though finding trustworthy peers may again be a prohibitive cost for small participants.
Anonymous Gossiping. Established protocols like Dandelion [11, 25] and Tor [22] have been considered in the past in the context of Ethereum [9, 42, 60]. The core idea of Dandelion is to first propagate messages along a path of a single node (the stem phase), before releasing a message (the fluff phase). This comes at a significant latency cost and is deemed incompatible with the economic incentives of publishing messages quickly before [9].
@@ 9.2 Defending Against DoS
The implications we outline in Section 8 contain two attack vectors that are based on DoS. A (partial) mitigation could thus consist of defending against DoS attacks. We consider this defense independently, as it is typically orthogonal to anonymity.
@@
Giuliari et al. [35] present a list of traditional defense strategies, such as IP-based filtering, cloud-based DoS protection, overprovisioning, and VPNs. The authors argue that none of these are satisfactory for consensus algorithms in a blockchain setting. For Ethereum specifically, these concerns are even more insurmountable, due to (i) Ethereum’s prohibitively large network size (see Appendix B) and (ii) the comparatively lower assumed economic power and hardware specifications of solo stakers. Giuliari et al. [35] instead propose a defense based on source authentication and rate limiting. While promising, source authentication represents a 180-degree shift from the source anonymity property required for danksharding, and would thus require significant changes.
@@ 10 Related Work
Network Measurements and Attacks. While many measurements have been done on Ethereum and other cryptocurrency networks, most relevant for our work are those that measure gossip properties of these networks, observed miner centralization in Bitcoin and Ethereum [20, 34, 48, 43]. Measurements of validator centralization at a network-level have been limited. In a previous protocol version of Ethereum [28], nodes in the network would subscribe to one (uniform) random attestation backbone per validator. Thus, previously, one could estimate lower bounds for how many validators were on a node. This was done in [13] where the authors estimated an average of 12K nodes in the network with 5-6K running at least one validator.
@@
Deanonymization attacks for cryptocurrencies have primarily taken advantage of network timing information, e.g., by showing the possibility of tracking the source of transaction data to the peer first to distribute the transaction in the Bitcoin [26, 8] and it’s Lightning Network [59], as well as the Ethereum network [61, 17]. Gossip protocols to prevent such deanonymization attacks have also been proposed, such as the family of Dandelion protocols [11, 25]. We highlight that the aforementioned timing-based deanonymization attacks require a connection to a large portion of the network to be sure of the originating source of messages for each address. Our attack, in contrast, takes advantage of protocol-level details that can be run from a single node to any number of its peers. The danger of the potential deanonymization of validators has been a topic of discussion on Ethereum for many years. However, to the best of our knowledge, apart from a meta-data-related technique [58, 40], no concrete deanonymization methodologies have been proposed.
@@
There is also a broader history of network-layer attacks in cryptocurrency networks including Eclipsing [38, 49, 39], broader network-partitioning attacks [62, 63], and network routing manipulations such as BGP-Hijacking [4] and its prevention [3] and miner-pool networks routing attacks [64]. Protocol-level details have also been utilized to infer peering relationships, such as the structure of transaction messages and their broadcast behavior in the Bitcoin [54, 21, 52] and Ethereum [46] networks, and peer discovery messages in the Monero network [18].
@@
Gossip Networks. GossipSub was introduced in 2020 [65]. An improved version (v.1.1) has been introduced a year later [66]. We are not the first to present a weakness of GossipSub. Kumar et al. for instance present a formal method analysis of GossipSub [44] and show that the score function is fair, albeit exploitable. In a companion paper [45], they go on to synthesize and simulate attacks on GossipSub.
@@ 11 Outlook
There are several ways to strengthen the attack and deanonymize a larger portion of the network. Our Rainbow nodes ran with closed ports, limiting connections to the reachable part of the network. Running them with open ports would expand our connections and likely deanonymize more validators. Additionally, we only consider the first message for each attestation, but including duplicates would further improve deanonymization. Lastly, our method focuses solely on attestations, but other validator tasks (e.g., aggregation and sync committees) also leak information that could enhance deanonymization, potentially identifying aggregators before they self-reveal.
@@
Appendix A Ethical Considerations

### v2

@@ 8 Insights
FR SO VA ZH unique
FR 14,388 2,842 5,326 10,769 2,675
SO 2,842 11,185 5,855 9,221 691
VA 5,326 5,855 52,916 27,717 23,577
ZH 10,769 9,221 27,717 132,443 93,854
Table 6: Pairwise overlap of deanonymized validators across machines, as well as the number of unique validators per machine. With the exception of the ZH node, half of the validators deanonymized on a Rainbow node have also been deanonymized by at least one other Rainbow node.
@@ 8.1 Geographical Distribution
On a continent basis, 46.79% of the peers we deanonymize are located in Europe, 38.05% in North America, 9.95% in Asia, 4.16% in Oceania, 0.77% in South America, and 0.27% in Africa. In Figure 8(a), we show the number of peers on which we deanonymized validators by Rainbow node as well as overall per country for the top ten countries. We deanonymized the largest number of peers in the United States making up 33.03% of the peers. This is more than double the number of peers we deanonymize in the next two biggest countries. This is similar to the overall geographical distribution of peers in the network, where the United States and Germany make up the biggest proportions. Additionally, we also observe a geographical bias of our Rainbow nodes in terms of the peers on which they deanonymize. Most notably, the VA Rainbow node deanonymizes more peers in the United States than the ZH Rainbow node. Similarly, the ZH node deanonymizes more peers in most European countries in comparison to the VA node.
@@
In Figure 8(b) we further provide insights regarding the number of validators deanonymized by each Rainbow node per country. We start by highlighting the most striking difference to the previous figure: most validators we locate are hosted in the Netherlands (i.e., 12.71%), while the United States only comes in fourth place. Note that a difference in the figures is expected given a non-uniform distribution of validators across machines. In general, we observe that there is a larger proportion of validators in Europe as opposed to peers with validators. In particular, 70.90% of the validators we locate are in Europe, 12.48% are in North America, 11.46% are in Asia, 5.08% are in Oceania, 0.05% are in Africa and 0.03% are in South America. Additionally, we again notice the same geographical bias, i.e., the SO node’s high relative proportion of deanonymizations in Australia and South Korea.
@@
We further repeat the preceding analysis for organizations instead of countries in Figure 9. The largest organization in terms of the number of peers is Comcast a residential ISP with 5.97% and the following three (Hetzner, OVH, and Amazon) are cloud providers (see Figure 9(a)). Around half of the peers we deanonymize are hosted through cloud providers, whereas the other half run through residential ISPs, i.e., home stakers.
@@
There is a large shift in these figures when we turn from peers (see Figure 9(a)) to validators (see Figure 9(b)). For one, eight out of the ten largest organizations are cloud providers and all of the biggest seven are cloud providers. We locate the largest number of validators in Amazon data centers, i.e., 19.07%. Surprisingly, the ZH node deanonymizes by far the largest proportion of validators in Amazon data centers even though it is the only Rainbow node not run from an Amazon data center. Finally, we highlight that around 90% of the validators are run through cloud providers, with the other 10% belonging to residential ISPs.
8.2 Validator Distribution
For those peers on which we deanonymize validators, we proceed by analyzing how many validators they host (see Figure 10). There is only a small number of validators located on most peers: a single validator on around 27% of our peers, with more than half of our peers having no more than four validators. In fact, we see that only 11% of peers have more than 100 validators. We further note that an in-detail inspection of Figure 10 reveals that for those peers that host many validators (i.e., more than 100), there is a bias towards a “round” number of validators to be hosted on a peer. This is evident by increases in the cdf at 100, 250, 300, 350, 400, 500, and 1,000 validators per peer. The gray vertical lines indicate these. For one, this indicates that large organizations running multiple validators on a machine tend to divide validators across nodes in round numbers as one would expect, e.g. an organization controlling 10,000 validators might divide them across 10 nodes and run 1,000 validators on each node. Additionally, this observation also indicates that our deanonymization technique locates all validators hosted on a machine as it is less likely for these “round” numbers to be a coincidence.
@@ 8.3 Staking Pools
In Figure 11, we plot a cdf of the number of validators hosted per peer for these five staking pools. Recall in our earlier analysis (see Figure 10) the observation that node operators of staking pools tend to run a “round” number of validators on the same machine. We observe these same increases in Figure 11. Here we observe an average of 709 validators on a given peer, with the larger validator set containing 19,390 validators. We highlight the security concern this raises: though there are a million validators, just hundreds of these peers going offline can stall Ethereum as more than two-thirds of the validators must be online to guarantee liveness.
@@
In Figure 12 we examine the distribution of validators in large liquid staking services in greater detail. The histogram shows the number of validators hosted by peers from the top five staking pools (four of which are deanonymized). While some peers host fewer than 100 validators, most peers in these pools host over 100, highlighting the economic benefits of staking pools. The most validators we locate on a single peer is 19,390.
@@
8.4 Implications on Decentralization
@@
Decentralization of the P2P network as well as Ethereum’s consensus layer is essential to prevent one party from maliciously overtaking the network. The decentralization of Ethereum’s PoS consensus layer is a topic of frequent discussion, given that the largest staking pool Lido controls nearly a third of the staking power – a critical safety threshold in Ethereum PoS. To combat the criticism, Lido proclaims that their validator set is operated by multiple independent node operators and thus they should all be seen as independent entities. While it is unclear whether this characterization is accurate [38], our insights into the P2P network reveal a further entanglement between different staking pools. As we outline in the previous section, not only do node operators run validators for different staking pools but we even find that they run them from the same nodes. This shows that staking pools are not always independent from each other, e.g., whenever such a node goes offline, validators from both protocols go offline. Similarly, we find that some peers host thousands of validators. Our insights gained from the deanonymization raise questions regarding the decentralization and resilience of the Ethereum consensus layer.
@@ 9.1 Providing Anonymity
We begin by discussing techniques that enhance anonymity within the P2P network. Before delving into the details, it is important to note that the current max effective balance for a validator is 32 ETH, which results in many redundant validators being run and controlled by the same entity. However, in the upcoming Pectra hardfork [65], the max effective balance is set to increase to 2048 ETH [62]. This change will allow entities to consolidate up to 64 validators into a single entity. Consequently, the next hard fork has the potential to significantly reduce the number of attestations propagated through the P2P network. This reduction allows for the following approaches that slightly increase the number of P2P messages to preserve anonymity. Importantly, with such an increase in P2P messages, the reductions in signature verification (i.e., one per validator per attestation) remain preserved.
@@
Solutions based on adding more nodes do not tackle the root cause of the issue (likely just increasing the complexity of deanonymization) and almost universally lead to higher message complexity. Furthermore, they also result in larger operating costs, that are especially difficult to bear for solo stakers, thus going against the ethos of Ethereum.
Private Peering Agreements. Out of the box, the Lighthouse and Prysm Ethereum clients provide the capacity to perform private peering agreements – a set of trusted peers can be defined who function as additional relays for gossip messages. This functionality is meant for increased performance and reliability, however, it also makes it harder to uniquely map validators to one IP. To be precise, through our presented deanonymization methodology, we could still determine the set of validators and the set of nodes that are part of the peering agreement, but not which validator is hosted by which node. Thus, a private peering agreement can provide k-anonymity [^20] In this context, a node is said to be k-anonymous if it cannot be distinguished from k-1 other nodes. As a result, it would be necessary to target the $k$ peers in the peering agreement during an attack. This mitigation comes with two main caveats: First, $k$ has to be chosen large enough to make DoS attacks prohibitively costly. Second, while peering agreements can be established off-band, finding trustworthy peers may again be a prohibitive cost for small participants, with the introduction of such complexities potentially contributing to centralization. Based on our results presented in Section 8, 50% of nodes hosting validators host less than four. Nonetheless, private peering agreements can at least mitigate against the threat of network-wide disruptions (as outlined in Section 3). The positive effect of such a mitigation can be seen today in validators who make use of service providers at the gossip layer (see Section 7).
Anonymous Gossiping. Established protocols like Dandelion [12, 29] and Tor [23] have been considered in the past in the context of Ethereum [10, 46, 70]. The core idea of Dandelion is to first propagate messages along a path of a single node (the stem phase), before releasing a message (the fluff phase). This comes at a significant latency cost and is deemed incompatible with the economic incentives of publishing messages quickly [10].
@@ 9.2 Defending Against DoS
Both main threats outlined in Section 3 can be exercised by employing DoS. A (partial) mitigation could thus consist of defending against DoS attacks. We consider this defense independently, as it is typically orthogonal to anonymity.
@@
Giuliari et al. [37] present a list of traditional defense strategies, such as IP-based filtering, cloud-based DoS protection, overprovisioning, and VPNs. The authors argue that none of these are satisfactory for consensus algorithms in a blockchain setting. For Ethereum specifically, these concerns are even more insurmountable, due to (i) Ethereum’s prohibitively large network size (see Section 5.2) and (ii) the comparatively lower assumed economic power and hardware specifications of solo stakers. Giuliari et al. [37] instead propose a defense based on source authentication and rate limiting. While promising, source authentication represents a shift for Ethereum, and would thus require significant changes.
@@
Distributed Validator Technology (DVT). By splitting the private key of a validator into multiple shares, using DVT a validator can run from multiple clients and create signatures even if some clients are unavailable [32]. The resilience gained may also help validators create blocks even when experiencing a DoS attack.
@@ 10 Related Work
Network Measurements and Attacks. The measurements of Ethereum and other cryptocurrency networks most relevant for our work are those that measure gossip properties of these networks, observed miner centralization in Bitcoin and Ethereum [21, 36, 54, 47]. Measurements of validator centralization at a network-level have been limited. In a previous protocol version of Ethereum [34], nodes in the network would subscribe to one (uniform) random attestation backbone per validator. Thus, previously, one could estimate lower bounds for how many validators were on a node. This was done in [14] where the authors estimated an average of 12K nodes in the network with 5-6K running at least one validator.
A large body of literature explores deanonymization attacks in the context of anonymity-preserving distributed systems, such as Crowds and Tor [6]. Foundational models [79, 31] demonstrate how repeated path observations degrade anonymity over time, exploiting statistical biases towards identifying initiators (so-called predecessor attacks). Crowds’ probabilistic routing amplifies this vulnerability compared to the fixed circuits used in Tor [64]. Other seminal works focus on traffic correlation [59], timing attacks [71], and learning-based attacks [60]. In comparison, our work exploits the performance improvements introduced in GossipSub, making our deanonymization attack more effective.
@@
For cryptocurrencies specifically, deanonymization attacks have primarily taken advantage of network timing information, e.g., by showing the possibility of tracking the source of transaction data to the peer first to distribute the transaction in the Bitcoin [30, 8, 9] and its Lightning Network [69], as well as the Ethereum network [72, 17]. Gossip protocols to prevent such deanonymization attacks have also been proposed, such as the family of Dandelion protocols [12, 29]. We highlight that the aforementioned timing-based deanonymization attacks require a connection to a large portion of the network to be sure of the originating source of messages for each address. Our attack, in contrast, takes advantage of protocol-level details that can be run from a single node to any number of its peers. The danger of the potential deanonymization of validators has been a topic of discussion on Ethereum for many years. However, to the best of our knowledge, apart from a meta-data-related technique [68, 44], no concrete deanonymization methodologies have been proposed.
@@
There is also a broader history of network-layer attacks in cryptocurrency networks including Eclipsing [40, 55, 42], network-partitioning [73, 74], and routing manipulations such as BGP-Hijacking [3] and its prevention [2] and miner-pool networks routing attacks [75]. Protocol-level details have also been utilized to infer peering relationships, such as the structure of transaction messages and their broadcast behavior in the Bitcoin [61, 22, 58] and Ethereum [51] networks, and peer discovery messages in the Monero network [18].
@@
Gossip Networks. GossipSub was introduced in 2020 [76]. An improved version (v.1.1) has been introduced a year later [77]. We are not the first to present a weakness of GossipSub. Kumar et al. [50] for instance present a formal method analysis of GossipSub and show that the score function is fair, albeit exploitable. In a companion paper [49], they go on to synthesize and simulate attacks on GossipSub.
@@ 11 Outlook
There are several ways to strengthen the attack and deanonymize a larger portion of the network. Our Rainbow nodes ran with closed ports, limiting connections to the reachable part of the network. Running them with open ports would expand connections and likely deanonymize more validators. Additionally, we only consider the first occurence of each attestation, but including duplicates would further improve deanonymization. Lastly, our method focuses solely on attestations, but other validator tasks (e.g., participation in aggregation and sync committees) also leak information. This information could not only enhance deanonymization, but potentially also identify aggregators before they self-reveal, opening up additional attack vectors.
@@
While this work focuses on Ethereum, it is important to emphasize that the foundational elements enabling the deanonymization attack are inherent to its use of the GossipSub [53] protocol. Developers planning to integrate GossipSub into their systems should carefully consider the privacy implications associated with its subnet overlay feature.
12 Ethics Considerations